OpenStudy (anonymous):

Of the following four numbers, which is the greatest? Select the best answer from the choices provided. Select one of the options below as your answer: A. 0.06 B. 0.6 C. 0.006 D. 0.66

OpenStudy (anonymous):

If you can't judge from given numbers, first convert all to like decimal nos 0.060 = 60/1000 0.600 = 600/1000 0.006 = 6/1000 0.660 = 660/1000 Now can you say which one is the greatest???

OpenStudy (anonymous):

Yeah its the one that has the most in the tenths and hundreths place^^^